hockey teams receive 1 point when they tie, and 2 points when they win. One season a team won with 64 points, they won 11 more games than they tied, how many wins and how many ties did they have?
Answer by checkley77(12844)

w=t+11
2w+1t=64
2(t+11)+t=64
2t+22+t=64
3t=64-22
3t=42
t=42/3
t=14 number of games tied.
w=14+11=25 games won.
Proof:
2*25+14*1=64
50+14=64
64=64